# CEERS: Forging the First Dust Grains in the Universe?

### v1

@@ Abstract
Context.
Aims.
Methods. This paper investigates the coevolution of metals and dust for 173 galaxies at observed with JWST/NIRSpec in the CEERS project. We focus on galaxies with extremely low dust attenuation to understand the physical mechanisms at play.
Results. We developed a new version of the CIGALE code that integrates spectroscopic and photometric data. By statistically comparing observations with modeled spectra, we derive physical parameters to constrain these mechanisms.
Conclusions. Our analysis reveals a population of 49 extremely low dust attenuation galaxies (GELDAs), consistent with within and . The stacked spectrum of GELDAs shows a very blue UV slope and a Balmer decrement H/H, consistent with no dust and Case B recombination with minimal underlying absorption. Notably, GELDAs are more prevalent at (83.3%) than at lower redshifts (26.3%), suggesting they could dominate in the early Universe. Using a far-infrared dust spectrum from the ALPINE sample, we study vs. trends. These exhibit upper and lower sequences connected by transitional galaxies. Our comparison with models indicates a critical transition around , from dust dominated by stellar sources (SNe and AGB stars) to dust growth via gas accretion. This corresponds to a metallicity of (), aligning with the point where ISM dust growth matches stellar dust production. The sample has a high gas fraction (), with no significant gas expulsion, and high surface gas densities. This leads to low star formation efficiencies compared to sub-millimeter galaxies. GELDAs may help explain the observed excess of bright galaxies at .
@@
In the last decade, we have detected dusty galaxies at z4 (Vieira et al. 2013; Hezaveh et al. 2015; Watson et al. 2015; Laporte et al. 2017; Fudamoto et al. 2017; Strandet et al. 2017; Tamura et al. 2019; Sommovigo et al. 2022a; Algera et al. 2024b, a; Witstok et al. 2023; Zavala et al. 2023; Valentino et al. 2024) with large dust masses (Pozzi et al. 2021; Akins et al. 2023) that cannot be explained by models. SNae and AGB stars could scarcely be at the origin of such a large dust mass, especially if we account for the reverse shock resulting from the expanding SN blast wave in the ISM (Leśniewska & Michałowski 2019). To explain such a large dust mass, we could assume an important role of dust mass growth in the ISM. However, this dust growth is regulated by some critical metallicity of the ISM (Inoue 2011; Asano et al. 2013; Zhukovska 2014; Feldmann 2015; Popping et al. 2017; Hou et al. 2019; Li et al. 2019; Graziani et al. 2020; Triani et al. 2020; Parente et al. 2022; Choban et al. 2024) to which the gas-dust accretion is turned on which, in turn, sets the transition from SNae and AGB formed stardust to accretion-dominated dust mass. Dust formation models, semi-analytical (Popping et al. 2017; Vijayan et al. 2019; Triani et al. 2020; Dayal et al. 2022; Mauerhofer & Dayal 2023) and cosmological (Graziani et al. 2020; Esmerian & Gnedin 2024; Lewis et al. 2023; Di Cesare et al. 2023; Lower et al. 2023; Choban et al. 2024) predict such a transition from galaxies only containing stardust created by SNae and AGB stars to galaxies where grain growth by accretion of metals in interstellar clouds becomes dominant. However, this transition has not yet been observationally confirmed in the high-redshift Universe.
Ferrara et al. (2022) suggest that dust could have been efficiently ejected during the very first phases of galaxy build-up. This hypothesis is supported by the fact that high-redshift galaxies are extremely bursty and very small (a few hundred parsecs (e.g. Sun et al. 2023; Choban et al. 2024). In this case, we would only detect the almost dust-free objects that remain. Both the star/ISM dust and the dust ejections would produce the same low dust attenuations and thus bright UV luminosities. Other discriminant observables such as the gas mass fraction and the metallicity should be explored to study this possible degeneracy. For example, one scenario could have a lot of gas (so something like CII could be detected) and no dust, while the other would have no gas nor dust.
@@ 3.2 The mass of metals and gas of the galaxy sample
We also need to add the contribution from the atomic gas Matomgas to Mmolgas The atomic-to-molecular mass ratio Matomgas/Mmolgas is estimated by Chowdhury et al. (2022) for star-forming galaxies at z0, z1.0, and z1.3 with values in the range 42 for galaxies with Mstar1010 M⊙. To account for galaxies with 1010 M⊙ in their statistics, they assume that the ratios Matomgas to Mmolgas are systematically higher by a factor of about 5 for all galaxies with Mstar M⊙. In this case, the values obtained would increase Mmolgas at z 1.3 by a factor of approximately 2, giving a ratio Matomgas/Mmolgas = 2.5 for the highest redshifts. At higher redshifts (0.01 z 6.4), there is no significant redshift evolution of the Matomgas/Mmolgas ratio (Messias et al. 2024), which is about 1-3. At z=8.496, the gas and stellar contents of a metal-poor galaxy are studied with JWST and ALMA (Heintz et al. 2023a). From this analysis, they infer Mmolgas = (3.0-5.0) 108 M⊙. corresponding to 40% 10% of Mgas for their object, which leads to Matomgas/Mmolgas = 1.5. Given the redshift of our objects, we will assume in this paper Matomgas/Mmolgas=2.0.
@@ 3.3 Dust masses
The far-IR information for this sample is limited and we cannot directly derive any information on the dust emission SED shape. However, the ALMA-ALPINE sample e.g. Béthermin et al. 2020; Pozzi et al. 2021; Sommovigo et al. 2022a and the ALMA-REBELS samples (e.g. Inami et al. 2022; Sommovigo et al. 2022b; Algera et al. 2024b and Algera et al. 2024a presents physical properties, e.g. stellar masses: (Mstar)10, redshifts: 4.5z6.2 similar to ours (Burgarella et al. 2022; Nanni et al. 2020). Due to the similarities of the samples, we will assume that we can make use of the same Draine et al. (2014) best-fit model (see Tab 3) identified in Burgarella et al. (2022); Nanni et al. (2020). We note that this ALMA-ALPINE model corresponds to a dust temperature Tdust=54.16.7 K, assuming an optically thin modified black-body, which is in good agreement with Sommovigo et al. (2022a) who found an average value =48 8 K and Mdust in the range (0.5-25.1) 107 M⊙ for ALPINE. For ALMA-REBELS, the median Tdust is in the 39-58 K range, and the median dust masses are estimated in the range (0.9-3.6) 107 M⊙ (Sommovigo et al. 2022b). Sommovigo et al. (2022b) also predict that dust masses can be produced by SNae alone for 85 % of the REBELS sample. We note that Algera et al. (2024b, a) found lower dust temperatures (Tdust=30 - 35 K) for two objects in the REBELS sample. Furthermore, Sommovigo et al. (2022a) predict that more metal-poor high-z galaxies could have warmer temperatures because of their smaller dust content, while the objectives studied in Algera et al. (2024b, a) are metal-rich. However, the dust model is assumed to be the same in our work for all of our objects. Globally modifying our model would cause an offset of all the dust masses, but not the observed relative difference between GELDAs and non-GELDAs.
@@
A quantization of the mean dust-to-metal, DTM=¡Mdust/MZ¿ shown in Fig. 13, gives DTMmean = 0.080, DTMmedian = 0.006 for GELDAs with a first quartile (25 %) Q1 = 0.002 and a third quartile (75 %) Q3 = 0.037, whereas for non GELDAs DTMmean = 0.654, DTMmedian 0.156 with Q1 = 0.085 and Q3 = 0.331. We observe a strong break in the DTM that could be interpreted (Inoue 2011; Asano et al. 2013; Zhukovska 2014; Feldmann 2015; Popping et al. 2017; Hou et al. 2019; Li et al. 2019; Graziani et al. 2020; Triani et al. 2020; Parente et al. 2022; Choban et al. 2024; Dubois et al. 2024) as hints that GELDAs have not started accretion growth of dust while non GELDAs are above the critical metallicity and have dust growth in the ISM. It is interesting to note that Rémy-Ruyer et al. (2014) observed a large scatter in the gas-to-dust mass ratio for a sample of 126 galaxies spanning a 2 dex range in metallicity. This scatter appears at 7.2 8.7 and is consistent with the dust growth in the ISM predicted by Asano et al. (2013) and other works cited in this paper. The objects are those appearing on the bottom left of Fig. 6 and Fig. 13, that is GELDAs. Observations show such a break in the DTM–metallicity relationship at around 0.1 Z⊙ (Rémy-Ruyer et al. 2014; De Vis et al. 2017) predicted by numerical models (e.g. Popping et al. 2017; Hou et al. 2019; Li et al. 2024; Parente et al. 2022. It is related to the transition in the leading mechanism of dust formation from stardust at low metallicity to ISM growth at high metallicity (Asano et al. 2013; Zhukovska 2014; Popping et al. 2017; Graziani et al. 2020; Li et al. 2019; Lewis et al. 2023.
@@ 4 Discussion
We now analyze the possible origins of GELDAs. To this aim, we first stack the spectra of the GELDAs. See Fig. 14 and line fluxes extracted from the stacked spectrum in Tab. 2. This spectrum is characteristic of star-forming galaxies, with a very blue UV slope = -2.451 0.066. This sample of GELDAs is compatible with no dust attenuation for Case B222from Groves et al. (2012): ”Case A and Case B. Case A assumes that an ionized nebula is optically thin to all Lyman emission lines, while Case B assumes that a nebula is optically thick to all Lyman lines greater than Ly, meaning these photons are absorbed and re-emitted as a combination of Ly and higher order lines, such as the Balmer lines. These two cases will lead to different intrinsic ratios for the Balmer lines, with variations of the same order as temperature effects. Although Case B is typically assumed for determining intrinsic ratios, in reality the ratio in typical H II regions lies between these two cases.”: H/H=2.932±0.660. To reach H/H=2.86 (no dust attenuation), we need to apply a correction to H for the underlying absorption of 2.5 %, a relatively low correction (Kashino et al. 2013; Reddy et al. 2015; Shivaei et al. 2020).

### v2

@@ Abstract
Aims. This paper investigates the coevolution of metals and dust for 173 galaxies at 4.0z11.4 spectroscopically observed by JWST/NIRSpec in the CEERS project. More specifically, we want to study and analyze the properties of a sample of galaxies that show an extremely low dust attenuation and try to understand the possible physical processes at play in these galaxies.
Methods. We develop a new version of the CIGALE code that accepts spectroscopic and photometric data. From a statistical comparison of the observations with the modeled spectra, we derive a set of physical parameters that allow us to constrain the above physical processes.
Results. Our analysis reveals a population of 49 extremely low dust attenuation galaxies (GELDAs) consistent with AFUV = 0.0 within 2 and Mstar 109M⊙. After stacking the spectra of the 49 GELDAs to increase the signal-to-noise ratio, we measure a very blue UV slope = -2.451 0.066, and a Balmer decrement H/H=2.9320.660 without underlying absorption and consistent with no dust attenuation and Case B assuming an underlying absorption of 2.5 %. Furthermore, the proportion of GELDA is much higher at z 8.8 (83.3 % of the total sample) than at z 8.8 (26.3 % of the total sample). This suggests that GELDAs become dominant in the early Universe. Assuming a prior far-infrared dust spectrum from the ALPINE sample, we perform an analysis of the properties of this galaxy population. The trends observed in the Mdust vs. Mstar diagram feature an upper and a lower sequence linked by objects that can be transitional. A comparison with models suggests that we might observe a critical transition at Mstar 108.5 M, corresponding to a critical metallicity of Zcrit = 12 + (O/H) 7.60 (i.e., Z/Z 0.1). At this point, galaxies transition from being dominated by stellar dust production (mainly from supernovae) to grain growth through gas–dust accretion in the ISM. The observational critical metallicity Zcrit derived in this paper is in good agreement with predictions from theoretical models for the onset of efficient grain growth. Furthermore, the mean gas mass fraction of our entire sample at 4.0z11.4 is very high: f. All of our galaxies, including GELDAs at all redshifts, contain a large amount of gas that was not expelled from the galaxies. Finally, the small size of the galaxies combined with the mass of gas lead to very high surface gas densities, which put our sample below high-redshift sub-millimeter galaxies, at relatively low star formation efficiency. The population of high-redshift GELDAs would provide us with a natural and inherent explanation for the origin of the apparent tension between observations and theoretical models in the number density of bright galaxies at z 9.
@@
In the last decade, we have detected dusty galaxies at z4 (Vieira et al. 2013; Hezaveh et al. 2015; Watson et al. 2015; Laporte et al. 2017; Fudamoto et al. 2017; Strandet et al. 2017; Tamura et al. 2019; Sommovigo et al. 2022a; Algera et al. 2024b, a; Witstok et al. 2023; Zavala et al. 2023; Valentino et al. 2024) with large dust masses (Pozzi et al. 2021; Akins et al. 2023) that cannot be explained by models. SNae and AGB stars could scarcely be at the origin of such a large dust mass, especially if we account for the reverse shock resulting from the expanding SN blast wave in the ISM (Leśniewska & Michałowski 2019). To explain such a large dust mass, we could assume an important role of dust mass growth in the ISM.
Models of dust formation, both semi-analytical (e.g., Popping et al., 2017; Vijayan et al., 2019; Triani et al., 2020; Dayal et al., 2022; Mauerhofer & Dayal, 2023) and cosmological (e.g., Graziani et al., 2020; Lewis et al., 2023; Di Cesare et al., 2023; Choban et al., 2024) predict a critical metallicity in the ISM above which grain growth via gas–dust accretion becomes efficient, marking the transition from dust production dominated by stellar sources (supernovae and AGB stars) to accretion-driven dust growth in the ISM. While this transition is supported by local-Universe observations (Rémy-Ruyer et al., 2014; De Vis et al., 2019), it remains undetected at high redshift.
Ferrara et al. (2022) suggest that dust could have been efficiently ejected during the very first phases of galaxy build-up. This hypothesis is supported by theoretical predictions that high-redshift galaxies are extremely bursty and very small (e.g. Sun et al. 2023; Choban et al. 2024). In this case, we would only detect the almost dust-free objects that remain. However, both star/ISM dust geometry and galactic dust ejections would produce the same low dust attenuations and thus bright UV luminosities. Other discriminant observables such as the gas mass fraction and the metallicity should be explored to study this possible degeneracy.
@@ 3.2 The mass of metals and gas of the galaxy sample
We also need to add the contribution from the atomic gas Matomgas to Mmolgas The atomic-to-molecular mass ratio Matomgas/Mmolgas is estimated by Chowdhury et al. (2022) for star-forming galaxies at z0, z1.0, and z1.3 with values in the range 42 for galaxies with Mstar1010 M⊙. To account for galaxies with 1010 M⊙ in their statistics, they assume that the ratios Matomgas to Mmolgas are systematically higher by a factor of about 5 for all galaxies with Mstar 1010M⊙. In this case, the values obtained would increase Mmolgas at z 1.3 by a factor of approximately 2, giving a ratio Matomgas/Mmolgas = 2.5 for the highest redshifts. At higher redshifts (0.01 z 6.4), there is no significant redshift evolution of the Matomgas/Mmolgas ratio (Messias et al. 2024), which is about 1-3. At z=8.496, the gas and stellar contents of a metal-poor galaxy are studied with JWST and ALMA (Heintz et al. 2023a). From this analysis, they infer Mmolgas = (3.0-5.0) 108 M⊙. corresponding to 40% 10% of Mgas for their object, which leads to Matomgas/Mmolgas = 1.5. Given the redshift of our objects, we will assume in this paper Matomgas/Mmolgas=2.0.
@@ 3.3 Dust masses
The far-IR information for this sample is limited and we cannot directly derive any information on the dust emission SED shape. However, the ALMA-ALPINE sample e.g. Béthermin et al. 2020; Pozzi et al. 2021; Sommovigo et al. 2022a and the ALMA-REBELS samples (e.g. Inami et al. 2022; Sommovigo et al. 2022b; Algera et al. 2024b and Algera et al. 2024a presents physical properties, e.g. stellar masses: (Mstar)10, redshifts: 4.5z7.7 similar to ours (Burgarella et al. 2022; Nanni et al. 2020). Due to the similarities of the samples, we will assume that we can make use of the same Draine et al. (2014) best-fit model (see Tab 3) identified in Burgarella et al. (2022); Nanni et al. (2020). We note that this ALMA-ALPINE model corresponds to a dust temperature Tdust=54.16.7 K, assuming an optically thin modified black-body, which is in good agreement with Sommovigo et al. (2022a) who found an average value =48 8 K and Mdust in the range (0.5-25.1) 107 M⊙ for ALPINE. For ALMA-REBELS, the median Tdust is in the 39-58 K range, and the median dust masses are estimated in the range (0.9-3.6) 107 M⊙ (Sommovigo et al. 2022b). Sommovigo et al. (2022b) also predict that dust masses can be produced by SNae alone for 85 % of the REBELS sample. We note that Algera et al. (2024b, a) found lower dust temperatures (Tdust=30 - 35 K) for two objects in the REBELS sample. Furthermore, Sommovigo et al. (2022a) predict that more metal-poor high-z galaxies could have warmer temperatures because of their smaller dust content, while the objectives studied in Algera et al. (2024b, a) are metal-rich. However, the dust model is assumed to be the same in our work for all of our objects. Globally modifying our model would cause an offset of all the dust masses, but not the observed relative difference between GELDAs and non-GELDAs.
@@
A quantization of the mean dust-to-metal, DTM=¡Mdust/MZ¿ shown in Fig. 13, gives DTMmean = 0.080, DTMmedian = 0.006 for GELDAs with a first quartile (25 %) Q1 = 0.002 and a third quartile (75 %) Q3 = 0.037, whereas for non GELDAs DTMmean = 0.654, DTMmedian 0.156 with Q1 = 0.085 and Q3 = 0.331. We observe a strong break in the DTM that could be interpreted (Inoue 2011; Asano et al. 2013; Zhukovska 2014; Feldmann 2015; Popping et al. 2017; Hou et al. 2019; Li et al. 2019; Graziani et al. 2020; Triani et al. 2020; Parente et al. 2022; Choban et al. 2024; Dubois et al. 2024) as hints that GELDAs have not started accretion growth of dust while non GELDAs are above the critical metallicity and have dust growth in the ISM. It is interesting to note that Rémy-Ruyer et al. (2014) observed a large scatter in the gas-to-dust mass ratio for a sample of 126 galaxies spanning a 2 dex range in metallicity. This scatter appears at 7.2 8.7 and is consistent with the dust growth in the ISM predicted by Asano et al. (2013) and other works cited in this paper. The objects are those appearing on the bottom left of Fig. 6 and Fig. 13, that is GELDAs.
@@ 4 Discussion
We now analyze the possible origins of GELDAs. To this aim, we first stack the spectra of the GELDAs. See Fig. 14 and line fluxes extracted from the stacked spectrum in Tab. 2. This spectrum is characteristic of star-forming galaxies, with a very blue UV slope = -2.451 0.066. This sample of GELDAs is compatible with no dust attenuation for Case B333from Groves et al. (2012): ”Case A and Case B. Case A assumes that an ionized nebula is optically thin to all Lyman emission lines, while Case B assumes that a nebula is optically thick to all Lyman lines greater than Ly, meaning these photons are absorbed and re-emitted as a combination of Ly and higher order lines, such as the Balmer lines. These two cases will lead to different intrinsic ratios for the Balmer lines, with variations of the same order as temperature effects. Although Case B is typically assumed for determining intrinsic ratios, in reality the ratio in typical H II regions lies between these two cases.”: H/H=2.932±0.660. To reach H/H=2.86 (no dust attenuation), we need to apply a correction to H for the underlying absorption of 2.5 %, a relatively low correction (Kashino et al. 2013; Reddy et al. 2015; Shivaei et al. 2020).